O'Brien Ltd.'s outstanding bonds have a $1,000 par value, and they mature in 20 years. Their nominal yield to maturity is 8.75%, they pay interest semiannually, and they sell at a price of $800. What is the bond's nominal (annual) coupon interest rate? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.
Answer #1

Information provided:

Par value= future value= $1,0000

Current price= present value= $800

Time= 20 years*2= 40 semi-annual periods

Yield to maturity= 8.75%/2= 4.3750%

The coupon rate is computed by calculating the amount of coupon payment.

Enter the below in a financial calculator to compute the coupon payment:

FV= 1,000

PV= -800

N= 40

I/Y= 4.3750

The value obtained is 33.0746.

The annual coupon payment is 33.0746*2= $66.1492.

Coupon rate= Annual coupon payment/ Par value

                        = $66.1492/ $1,000

                        = 0.0661*100

                        = 6.61.

In case of any query, kindly comment on the solution.